What is the probability of drawing three cards out of a standard deck of 52 cards, without replacement, and have all 3 cards turn up red?
Answer by rothauserc(4718)

There are 26 red cards in the deck of 52, therefore we have three probabilities to multiply, they are
(26/52) * (25/51) * (24/50) = (15600 / 132600) = .1176 = .12